OpenStudy (anonymous):

find the area of a triangle. (A=1/2 bh) triangle height=2x base=2x^2

OpenStudy (anonymous):

ok so you have \[A = \frac{b*h}{2}\] \[h = 2x\] \[b = 2x^2\] \[A = \frac{(2x^2)*(2x)}{2} = \frac{4x^3}{2} = ?\]

OpenStudy (anonymous):

hm i think i might have calculated wrong because the answer i got isnt right. i did the 4x^3 divided by 2 but didnt get the right answer.

OpenStudy (anonymous):

4/2 = 2 so you should get 2x^3